HUBBLE SPACE TELESCOPE DAILY REPORT #1991 PERIOD COVERED: 0000Z (UTC) 10/24/97 - 0000Z (UTC) 10/27/97 Daily Status Report as of 300/0000Z 1.0 ACTIVITIES SCHEDULED AND ACCOMPLISHED: 1.1 Completed Three Sets of STIS/CCD 7802 (Daily Darks to Update Acquisition Bad Pixel Table) The Space Telescope Imaging Spectrograph (CCD) was used to make calibration observations to obtain daily darks to update the acquisition bad pixel table. These darks are intended to identify hot pixels for update to the acquisition on-board FSW hot pixel table. Daily hot pixel lists can also be used for science data analysis. The first iteration of this proposal was affected by a guide star acquisition failure described in the previous report, and detailed in HSTAR 6216. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.2 Completed NIC/1/2/3 7726 (NICMOS One Orbit Parallel) The Near Infrared Camera (Cameras One, Two and Three) was used to make one orbit parallel calibration observations. This proposal was affected by a guide star acquisition failure described in the previous report, and detailed in HSTAR 6216. 1.3 Completed Two Sets of STIS/CCD 7781 (STIS Non-Scripted Parallel Archive Proposal - Full Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one full orbit). The first iteration of this proposal was affected by a guide star acquisition failure described in the previous report, and detailed in HSTAR 6216. Otherwise, the observations were executed as scheduled, and no further anomalies were noted. 1.4 Completed Five Sets of WF/PC-2 7621 (WF/PC-2 Cycle 7 Supplemental Darks PT1/3) The WF/PC-2 was used to make supplemental dark calibration images. This set of calibrations is intended to obtain three dark frames every day to provide data for monitoring and characterizing the evolution of hot pixels. The observations were executed as scheduled, and no anomalies were noted. 1.5 Completed NIC/3 7806 (NICMOS GRISM Verification Test) The Near Infrared Camera (Camera Three) was used to perform a verification test by making observations of VY2-2. This is a validation of the operations of the two NICMOS grisms used in the HDF observations, and supply the initial level of calibration for spectroscopic observations. The observations were executed as scheduled, and no anomalies were noted. 1.6 Completed Nineteen Sets of WF/PC-2 7786 (Pure Parallel Archive Survey - Continuation) The WF/PC-2 was used to make observations as part of an archival survey program in which the images were taken in parallel with the prime science observations. These WF/PC-2 parallel observations were obtained in the ultraviolet in a search for interesting objects. The observations were completed as planned, and no problems were reported. 1.7 Completed FGS 6876 (Parallax of RR-Lyrae) The FGSs were used to make astrometric observations of the variable star RR-Lyrae. RR Lyrae is the prototype of a class of stars that provide a fundamental stepping stone for the astronomical distance scale. Its trigonometric parallax is presently known to be around 0.0042 +/- 0.0019 arcsec for a distance modulus of 6.86 +/- 0.6 mag. It is the only star of this class within parallax range of HST. The observations were completed as planned, and no problems were reported. 1.8 Completed Two Sets of NIC/2 7498 (NICMOS Imaging Survey of Distant Radio Galaxies. Ellipticals at Z > 2 ?) The Near Infrared Camera (Camera Two) was used to make high resolution and high signal-to-noise ratio imaging survey observations of the radio galaxies MRC2104-24 and USS1707+10. The existence and photometric properties of massive elliptical galaxies at high redshift provide a powerful probe of galaxy evolution. The observations were executed as scheduled, and no anomalies were noted. 1.9 Completed Twenty Three Sets of STIS/CCD 7783 (STIS Parallel Science) The Space Telescope Imaging Spectrograph (CCD) was used to make observations taken in parallel with the prime science. The observations were completed, and no problems were reported. 1.10 Completed Two Sets of WF/PC-2 6673 (Black Holes and Gas Disks in a Complete Sample of Radio Loud UGC Ellipticals) The WF/PC-2 was used to make observations of radio load, elliptical galaxy UGC-00408 in order to search for possible black holes and gas disks. Hydrogen-Alpha and broad band observations were taken to determine the luminosity profiles and stellar populations (on the tens of parsec scale), and to detect nuclear regions or disks of gas and dust. The observations were completed as planned, and no anomalies were reported. 1.11 Completed Four Sets of NIC/1/2/3 7729 (NICMOS Short One Orbit Parallel) The Near Infrared Camera (Cameras One, Two, and Three) was used to obtain "JHK" imaging of sky regions while another instrument was prime. These observations were part of the short NICMOS one-orbit parallel program. The observations were executed as scheduled, and no anomalies were noted. 1.12 Completed Two Sets of STIS/CCD 7782 (STIS Non-Scripted Parallel Archive Proposal - Half Orbit) The Space Telescope Imaging Spectrograph (CCD) was used for a non-scripted parallel archive proposal (one-half orbit). The observations were executed as scheduled, and no anomalies were noted. 1.13 Completed STIS/CCD 7642 (CCD Red Light PSF Halo) The Space Telescope Imaging Spectrograph (CCD) was used to measure the PSF of the red star G158-100 with the CCD at red wavelengths, to characterize the internal scattering of long-wavelength light within the CCD itself. The observations were completed as planned, with no reported problems. 1.14 Completed Six Sets of WF/PC-2 7625 (WF/PC-2 Cycle 7 Earth Flats) The WF/PC-2 was used to make a series of Earth flat calibration images. These flat fields will allow mapping of the OTA illumination pattern and will be used in conjunction with previous internal and external flats to generate new pipeline superflats. The observations were completed as planned, and no anomalies were reported. 1.15 Completed NIC/2 7331 (Near-IR Properties of the Bulges of Spiral Galaxies) The Near Infrared Camera (Camera Two) was used to make infrared observations of the bulge in spiral galaxy NGC-2339. Near-IR data at HST resolution are essential for correcting for the effect of dust and will enable us to derive robust nuclear properties for the sample. The observations were executed as scheduled, and no anomalies were noted. 1.16 Completed Two Sets of NIC/1/2/3 7714 (Observations of the Rest-Frame Optical Morphology of a Statistically Significant Sample of Normal Galaxies at Redshift z~3.) The Near Infrared Camera (Cameras One, Two and Three) was used to make observations of galaxy 003-2603. It is planned to use 30 orbits with NICMOS to study the rest-frame optical morphology of star-forming galaxies at redshift z~3 selected with the Lyman-limit technique. The observations were executed as scheduled, and no anomalies were noted. 1.17 Completed WF/PC-2 6853 (Emission Features Near Eclipsed Io) The WF/PC-2 was used to make observations of Jupiter's moon Io during an eclipse of the moon. The high spatial resolution of the WF/PC-2 will allow an extension of the ground based observations of interactions between Io and the Jovian plasma torus. The observations were completed as scheduled, and no anomalies were reported. 1.18 Completed COSTAR/U 6895 (FOC Cycle 6 PSF Monitoring) The Faint Object Camera and COSTAR were used to make observations of object 01-001 (aka, white dwarf star BPM16274) in order to check out the Point Spread Function (PSF) and focus of the system. The observations were executed as scheduled, and no problems were reported. 1.19 Completed NIC/1/2 7443 (Mineralogic Mapping of Asteroid 4 Vesta) The Near Infrared Camera (Cameras One and Two) was used to make mineralogic mapping of asteroid 4 Vesta. The 1997 opposition of asteroid 4 Vesta offers an important opportunity to map its diverse mineralogy with uniquely high contrast and to investigate more thoroughly the newly discovered major topographic feature at the South pole. The observations were completed as planned, and no anomalies were reported. 1.20 Completed NIC/2/3 7424 (NICMOS Imaging of QSO Absorption-- Selected Galaxies at z > 1) The Near-Infrared Camera (Camera Two and Three) was used to make observations of QSO/galaxy PG1329+412. The images will also be used to connect the absorption line properties to the galaxy morphological properties, inclination angles, and orientations in order to continue to explore the connection between the absorbing gas and the galaxies. The observations were completed as scheduled, and no anomalies were reported. 1.21 Completed STIS/CCD 7711 (STIS CCD G750L Fringing Flats.) The Space Telescope Imaging Spectrograph (CCD) was used to investigate and monitor the STIS CCD flat-field fringing in the G750L grating at it's nominal wavelength setting {7751A}. WAVECALs were also taken at the same gain and wavelength setting. Typical S/N = 100 for each integration, and GAIN=4 on each. The observations were executed as scheduled, and no anomalies were noted. 1.22 Completed NIC/1/2/3 7158 (NICMOS Detector Linearity and Saturation) The Near Infrared Camera (Cameras One, Two and Three) was used to test the detector linearity and saturation of the camera. Some evidence has been found in SMOV and ERO data of changes in saturation level for NICMOS detectors since SLTV. This SMOV proposal aims to investigate this by obtaining linearity curves and saturation levels for all three NICMOS detectors, using a technique which avoids errors due to persistence. The observations were completed as planned, and no anomalies were reported. 1.23 Completed STIS/CCD 7721 (Slit Throughputs) The Space Telescope Imaging Spectrograph (CCD) was used to make observations of standard source FEUGE110 in order to determine the relative throughput vs. wavelength for a variety of commonly used slits. The observations were executed as scheduled, and no anomalies were noted. 1.24 Completed NIC/2 7241 (Near-IR Spectral Reflectance of the Jupiter Ring) The Near Infrared Camera (Camera Two) was used to make near infrared observations of Jupiter's ring in order to study its spectral reflectance. The near-infrared spectral reflectance of the Jupiter dust ring is poorly known because of problems with scattered light from the planet. The loss of lock mentioned in 2.1 occurred near the beginning of this proposal. Otherwise, the observations were completed as planned, and no further problems were reported. 1.25 Completed NIC/2 7219 (Ultraluminous Infrared Galaxies) The Near Infrared Camera (Camera Two) was used to make observations of the ultraluminous infrared galaxy IRAS17208. The luminous IRAS Galaxies are probably powered by nuclear starbursts and/or central AGNs; many of these systems exhibit evidence of recent galactic interaction or merging. The observations were completed as planned, and no problems were reported. 2.0 FLIGHT OPERATIONS SUMMARY: 2.1 Guide Star Acquisitions: Scheduled Acquisitions : 20 Successful : 18 Scheduled Reacquisitions : 26 Successful : 25 Per HSTAR 6219, there was a guide star loss of lock at 299/093708Z. The proposal described in 1.24 was affected. As explained in HSTAR 6221, the guide star acquisition scheduled at 300/0634Z failed and introduced a high roll error which caused the two following FHST updates to fail. This error also caused the acquisiton at 300/0815Z to fail as well as the subsequent re-acquisition at 0942Z. A realtime map was performed and an attitude reference update was uplinked at 300/104602Z. An FGS, two NICMOS , and a WF/PC proposal were affected. These will be identified in tomorrow's report. 2.2 FHST Updates: Scheduled : 47 Successful : 44 The updates at 299/0937Z failed. The final two updates failes are addressed in comments in 2.1. 2.3 Operations Notes: As mentioned in the last report, the Battery 1 discharge was started at 296/2229Z. The discharge was terminated nominally at 15 volts by a special hardware limit test at 297/1442Z. There is an estimated capacity of 78.5 A-h down to 26.37 volts. The previous capacity of 83.7 A-h was measured in March 1997. Recharge was started at 297/1510Z. The charge acceptance behavior is typical, relative to the other batteries. The six battery configuration will be returned to later today. The STIS interrupt log was dumped at 298/2054Z. The STIS MAMA2 MCE reset outside any SAA boundaries at 298/230902Z. It was recovered in the SMS at 300/0307Z. This was the first time that this method had been utilized. No STIS proposals were affected. As described in HSTAR 6220, the STIS was suspended at 300/052926 while processing a relative time command request. Indications were that the sequence began with an illegal command. After consultation between SE and the STScI, a STIS memory dump was performed at 300/0742Z per ROP NS-10. A NSSC-1 memory dump was performed at 300/1054Z per ROP NS-2. Further recovery operations will be performed today. 3.0 SIGNIFICANT FORTHCOMING EVENTS: Continuation of normal science observations and calibrations. /DMH